deathsindragon:

solve for v 2v^2+5=-7v

Da:

hmm i dont know

jasmine241015:

you can use mathway to get that answer hope it helps

deathsindragon:

its blocked

supie:

ok lemme help u out a little ight creep

supie:

what do you think is the first step?

deathsindragon:

if u gonna call me a creep u ca get the hell off my question

supie:

ight bet

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jasmine241015 you can use mathway to get that answer hope it helps \(\color{#0cbb34}{\text{End of Quote}}\) Yup

deathsindragon:

like i said its blocked

angiesyl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie ight bet \(\color{#0cbb34}{\text{End of Quote}}\) im dedd

angiesyl:

anyways the answer is

jasmine241015:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @deathsindragon like i said its blocked \(\color{#0cbb34}{\text{End of Quote}}\) oh I didn't know that that's the only math app ik of

angiesyl:

Step 1: Subtract -7v from both sides. 2v2+5−−7v=−7v−−7v 2v2+7v+5=0 Step 2: Factor left side of equation. (2v+5)(v+1)=0 Step 3: Set factors equal to 0. 2v+5=0 or v+1=0 v= −5 2 or v=−1

angiesyl:

Answer: v= −5 2 or v=−1

angiesyl:

pretty sure dis it

jasmine241015:

deathsindragon:

thank u guys

Lilian9667:

Your Welcome